# Consensus formation in quality-sensitive interdependent agent systems

### v1

@@ I Introduction
Many efforts have been devoted in the modelling of opinion dynamics in order to understand how a collective makes a decision among a set of options. Initial attempts devised models that encompassed simple imitation rules controlling the build up of an agreement between equivalent options, such as the voter model [9, 10] or the majority rule model [11]. This models, though, have been regarded as too unrealistic in their basic formulation, and many more complex models have been proposed in order to capture a richer spectrum of social behaviors, such as partisanship, heterogeneity or personalized information [12, 13, 14, 15].
@@
In this paper, we delve deeper into studying the formation of consensus by leveraging the LES model. Our investigation focuses on the interplay between social feedback and individual exploration, which the model balances through a single parameter known as the “interdependence” parameter. The intensity of individual, independent exploration serves a dual purpose. On one hand, it is essential for the group to perceive all available options or to adapt in changing environments [26, 30]. On the other hand, excessive individual exploration can introduce noise that hampers the effectiveness of the aforementioned positive feedback loop. Agents tend to overlook the actual state of their peers and adopt opinions based on the outcome of their own exploration. Consequently, stronger social interaction is required to balance this effect. Recent work highlights the capability of decentralized systems to effectively balance noise with social interactions under various scenarios [31, 32, 33].
@@ II Modeling fully-connected scout-bee-like swarms
The first term of this equation represents the likelihood that a scout bee independently discovers site $\alpha$, while the second term quantifies the probability that the bee commits to state $\alpha$ by following the advice of its peers: $f_{\alpha,t}$ denotes the fraction of agents promoting site $\alpha$ at time $t$. These terms are weighted by the interdependence parameter, $\lambda$. This parameter, ranging between 0 and 1, determines the extent to which bees depend on each other to reach their commitment decision. Vanishing interdependence ($\lambda\rightarrow 0$) signifies that bees will commit to a site solely based on independent exploration. Conversely, high interdependence ($\lambda\rightarrow 1$) means that bees will mostly consider their peers advertisement in order to make a commitment decision. Probabilities must be normalized, so Eq. (2) must satisfy $\sum_{\alpha=0}^{k}p_{\alpha,t}=1$. It’s important to note that we include the uncommitted state 00 in the normalization condition. Additionally, the normalization condition implies that $\sum_{\alpha=0}^{k}\pi_{\alpha}=1$, and consequently $\sum_{\alpha=1}^{k}\pi_{\alpha}\leq 1$.
The uncommitment transition probabilities ${r_{\alpha,t}}$ are inherently linked to the qualities of the sites. In the original LES model, transitions from committed to uncommitted states occur deterministically after a predetermined commitment time has elapsed. An agent commits to an option and advertises it for a fixed duration. To render the equations of motion mathematically tractable, T. Galla [29] replaced this deterministic process by a stochastic process defined by the following rates:
@@ II.1 Stationary deterministic solutions
Eq. (6) can be solved either numerically or rearranged into a $k+1$-th degree polynomial. In either approach, we find $k+1$ roots for $f_{0}$, some of which may lead to nonphysical solutions where $f_{0}^{*}>1$. A linear stability analysis is provided in Appendix A.1. In the specific case of $k=2$, assuming $q_{2}>q_{1}$ (or $r_{2}<r_{1}$), a single physical and stable solution $f^{*}_{j}$ $\forall j$ is obtained. An example of such a solution is depicted in Fig. 2A, with values $r_{2}=1/10$, $r_{1}=1/7$, $\pi_{1}=\pi_{2}=0.1$, $\lambda=0.6$ and $\mu=0$.
@@ II.1.2 Limit of fully interdependent bees
Although all these solutions are physically valid, only two of them (one in the case $\lambda=1$) are stable. Using linear stability analysis (LSA), we determine that the stability threshold between these two solutions is determined by the highest site quality, denoted as site $k$. When $\lambda<r_{k}$, the stable solution is the absorbing state $f_{0}=1$. Thus, in the absorbing state, the system remains fully uncommitted. Conversely, when $\lambda>r_{k}$, the stable solution is the one with the best site taking the lead, with $f_{\alpha}^{*}=(1-\frac{r_{k}}{\lambda})\delta_{\alpha,k}$, and the remaining proportion of uncommitted population being $f_{0}^{*}=\frac{r_{k}}{\lambda}$. In the special case where we set $\lambda=1$, the stability threshold vanishes because $\lambda=1\geq r_{k}$, and the only stable solution remaining is $f_{0}^{*}=r_{k}$, $f_{\alpha}^{*}=(1-r_{k})\delta_{\alpha,k}$. Details on the LSA analysis are provided in Sec. A.2.
In Fig. 2B,D, we present a summary of the findings for the binary problem with $k=2$. When $\lambda\geq r_{2}$, only the fixed point where the best option is imposed remains stable, while the one where the inferior option is imposed acts as a saddle node (stable on the simplex $f_{2}=0$). As $\lambda$ decreases, these points gradually move upward (reducing either $f_{1}$ or $f_{2}$) until they merge with the point $f_{0}=1$ at $\lambda=r_{\alpha}$. In the regime where $\lambda<r_{2}$, the only remaining stable fixed point is the absorbing state $f_{0}=1$ fixed point.
@@ II.1.3 Limit of equal-quality sites
Figure 3: Color-maps illustrating the stationary values of $f_{0}$, $f_{1}$, and $f_{2}$ in the parameter space $\pi-\lambda$ are presented in panels A, B, and C, respectively. The site qualities are held constant at $(q_{1},q_{2})=(7,10)$. Color gradients are used to indicate the quantitative values of $f_{\alpha}$ based on the color bars, while continuous lines represent iso-population levels in the parametric space.
@@ III Results
We will first concentrate on the specific case of a binary decision problem, a topic extensively studied in various general opinion dynamics models  [12, 11, 10, 35, 13, 36]. This scenario has also been the focus of other honeybee-inspired models [20, 21, 22, 23]. The generalization to larger values of $k$ is discussed later in Section III.6.
Figure 4: A,B: Color-maps illustrating the consensus parameters $Q_{2}$ and $Q_{2}^{\prime}$, as defined in Eqs. (11) and (12) respectively with $x=2$, across the parameter space $\lambda$-$\pi_{1,2}$. The black lines represent the crossover points $Q_{2}=0$ or $Q_{2}^{\prime}=0$. C: Time to reach the stationary state, $t_{ss}$, obtained from simulations with a swarm size of $N=5000$ bees. The fully interdependent limit ($\pi_{\alpha}=0$ or $\lambda=1$) is not included in the maps (see discussion in Sec.III.4). The site quality parameters are held constant at $(q_{1},q_{2})=(7,10)$ across all three maps. D,E: Crossover lines $\lambda^{*}(Q_{2},Q_{2}^{\prime}=0)$ as a function of $\pi_{1,2}$. Each curve represents a different value of $q_{1}$, while $q_{2}=10$ is held constant. F: Time to reach the stationary state, $t_{ss}$, plotted against the quality of the inferior option, $q_{1}$, while maintaining $q_{2}=10$. The plots depict three distinct values of interdependence, $\lambda$, with fixed discovery probabilities $\pi_{1}=\pi_{2}=0.1$. The markers are color-coded based on the consensus established at the stationary state. They are slightly jittered around the $q_{1}$ axis to allow for better visualization of the color code.
As detailed in the model description section, we consistently assume that site 2 holds the highest quality, denoted by $q_{2}>q_{1}$. Additionally, we assume that bees independently assess the qualities of the sites, represented by the model parameter $\mu=0$. This implies that the abandonment rates of the dances depend solely on the qualities of each site, defined as $r_{\alpha}=q_{0}/q_{\alpha}$. Previous research on this model has shown that, under these conditions, the swarm can identify the best site across a wide range of interdependence values, $\lambda$ [19, 29]. In this setup, we investigate the interplay between group communication, represented by the model parameter $\lambda$, and individual exploration success, captured by $\pi_{j}$. We explore scenarios where sites may vary in their likelihood of being discovered, denoted by $\pi_{1}\neq\pi_{2}$ (asymmetric scenario), or where they share the same discovery probability, denoted by $\pi_{1}=\pi_{2}$ (symmetric scenario).
@@ III.1 Symmetric discovery scenario
Fig. 4D illustrates the consensus crossover lines for different choices of $q_{1}$, while keeping $q_{2}=10$ fixed. When we increase the quality of the bad option, consequently increasing its advertisement time, the system requires a higher degree of interdependence to favor the higher quality option. This trend would reach a turning point when $q_{1}=q_{2}$. In this scenario, the system reaches a deadlock, where $f_{1}=f_{2}$. Conversely, when $q_{1}$ is reduced, the crossover line shifts to lower values of $\lambda^{*}$, potentially leading to a situation where interdependence is not necessary to achieve consensus. For instance, the crossover line for $q_{1}=5$ lies along the x-axis, with $\lambda^{*}=0$ for all $\pi$. In this case, the difference in advertisement time is sufficient to maintain a significant difference between $f_{2}$ and $f_{1}$, facilitating consensus even in the absence of interdependence. Furthermore, it’s notable that the consensus crossover lines terminate abruptly at the point $(\pi_{1,2},\lambda)=(0.0,0.1)$. As we demonstrated earlier in the fully interdependent limit ($\pi_{1,2}=0$), the system stays fully uncommitted until the threshold $\lambda=r_{2}=0.1$. Beyond this threshold, $f_{2}$ begins to take positive values while $f_{1}$ remains zero, resulting in trivial consensus.
Figure 5: Color-maps illustrating the stationary values of $f_{0}$, $f_{1}$ and $f_{2}$ are presented in panels A, B, and C, respectively, for the asymmetric discovery scenario in the parameter space $\pi_{2}-\lambda$. Here, the site quality parameters are fixed at $(q_{1},q_{2})=(7,10)$, and the discovery probability for site 1 is set to $\pi_{1}=0.25$. Color gradients indicate the quantitative values of $f_{\alpha}$ according to the color bars, while continuous lines represent iso-population levels in the parametric space.
@@
Figure 6: A: Color-map illustrating the stationary value of the consensus parameter $Q_{2}$ (as defined from Eq. (11) with $x=2$) in the parametric space $\lambda$-$\pi_{2}$. Site quality parameters are fixed at $(q_{1},q_{2})=(7,10)$, and the discovery probability for site 1 is set to $\pi_{1}=0.25$. The solid line indicates $Q_{2}=0$, while the dashed line indicates a simple majority, $Q_{1}=0$. B: Crossover lines $\lambda^{*}$ above which the consensus $Q_{2}$ becomes positive as the self-discovery probability $\pi_{2}$ is varied. Each curve represents a different value of site-1 quality $q_{1}$, while $q_{2}=10$ and $\pi_{1}=0.25$ are held constant. C: Crossover lines $\lambda^{*}$ above which $Q_{2}$ becomes positive as $\pi_{2}$ is varied. In this panel, qualities are held constant, $(q_{1},q_{2})=(7,10)$, and we illustrate the effect of different values of $\pi_{1}$. D: Time to reach the stationary state $t_{ss}$ in the state space ($\pi_{2}/\pi_{1},\lambda)$ obtained from simulations. Parameters $\pi_{1}=0.25$, $(q_{1},q_{2})=(7,10)$ and $N=5000$ are fixed. E: Time to reach the stationary state, $t_{ss}$, plotted against the quality of the inferior option, $q_{1}$, while maintaining $q_{2}=10$. Parameters $(\pi_{1},\pi_{2})=(0.25,0.05)$ and $N=5000$ are fixed. Color markers indicate the consensus value attained at the stationary state. Simulations start from uncommitted initial conditions.
@@
The trend observed is similar to that of consensus: interdependence tends to increase the decision time across all values of $\pi_{1,2}$, while the discovery probabilities speed up the process, reducing $t_{ss}$. This correlation between the established consensus value and the time to reach the stationary state suggests a speed-accuracy trade-off, commonly observed in biological systems [37] an in models of collective decision making [25, 36]. Achieving the best decision typically takes longer at higher values of $\lambda$, while processes are quicker below the consensus crossover line, where a less accurate consensus, such as a simple majority, can be reached much faster.
Similarly, increasing the value of the discovery probabilities, and thus the introduction of independent information, accelerates the decision process at the cost of a lower value of $Q^{\prime}$, which can be even negative, depending on $\lambda$. Only at the limits $\lambda\rightarrow 0$ or $\lambda\rightarrow 1$ $\pi_{1,2}$ has practically no effect: in the former, there’s no discussion mediated by interdependence, resulting in a quickly established yet inaccurate stationary state, while in the latter, social feedback dominates the process, with little impact from individual exploration.
It’s interesting to note the agreement between the stationary times discussed here and the relaxation times obtained from the linear stability analysis eigenvalues, as depicted in Fig. A1.

### v2

@@ I Introduction
Many efforts have been devoted in the modeling of opinion dynamics in order to understand how a collective makes a decision among a set of options. Initial attempts devised models that encompassed simple imitation rules controlling the build up of an agreement between equivalent options, such as the voter model [9, 10] or the majority rule model [11]. This models, though, have been regarded as too unrealistic in their basic formulation, and many more complex models have been proposed in order to capture a richer spectrum of social behaviors, such as partisanship, heterogeneity, personalized information or non linear interactions [12, 13, 14, 15, 16, 17].
@@
In this paper, we delve deeper into studying the formation of consensus by inspecting the LES model. Our investigation focuses on the interplay between social feedback and individual exploration, which the model balances through a single parameter known as the “interdependence” parameter. The intensity of individual, independent exploration serves a dual purpose. On one hand, it is essential for the group to perceive all available options or to adapt in changing environments [28, 32]. On the other hand, excessive individual exploration can introduce noise that hampers the effectiveness of the aforementioned positive feedback loop. Agents tend to overlook the actual state of their peers and adopt opinions based on the outcome of their own exploration. Consequently, stronger social interaction is required to balance this effect. Recent work highlights the capability of decentralized systems to effectively balance noise with social interactions under various scenarios [33, 34, 35].
@@ II Modeling fully-connected scout-bee-like swarms
The first term of this equation represents the likelihood that a scout bee independently discovers site $\alpha$, while the second term quantifies the probability that the bee commits to state $\alpha$ by following the advice of its peers: $f_{\alpha,t}$ denotes the fraction of agents promoting site $\alpha$ at time $t$. These terms are weighted by the interdependence parameter, $\lambda$. This parameter, ranging between 0 and 1, determines the extent to which bees depend on each other to reach their commitment decision. Vanishing interdependence ($\lambda\rightarrow 0$) signifies that bees will commit to a site solely based on independent exploration. Conversely, high interdependence ($\lambda\rightarrow 1$) means that bees will mostly consider their peers advertisement in order to make a commitment decision. Probabilities must be normalized, so Eq. (2) must satisfy $\sum_{\alpha=0}^{k}p_{\alpha,t}=1$. It’s important to note that we include the uncommitted state 00 in the normalization condition. Additionally, the normalization condition implies that $\sum_{\alpha=0}^{k}f_{\alpha}=1$, and consequently $\sum_{\alpha=1}^{k}\pi_{\alpha}\leq 1$.
The uncommitment transition probabilities ${r_{\alpha,t}}$ are inherently linked to the qualities of the sites. In the original LES model, transitions from committed to uncommitted states occur deterministically after a predetermined commitment time has elapsed. An agent commits to an option and advertises it for a fixed duration. To render the model’s equations mathematically tractable, T. Galla [31] replaced this deterministic process by a stochastic process defined by the following rates:
@@ II.1 Stationary deterministic solutions
Eq. (6) can be solved either numerically or rearranged into a $k+1$-th degree polynomial. In either approach, we find $k+1$ roots for $f_{0}$, some of which may lead to nonphysical solutions where $f_{0}^{*}>1$. A linear stability analysis is provided in Appendix A.1. In the specific case of $k=2$, assuming $q_{2}>q_{1}$ (or $r_{2}<r_{1}$), a single physical and stable solution $f^{*}_{\alpha}$$\forall\alpha$ is obtained. An example of such a solution is depicted in Fig. 2A, with values $r_{2}=1/10$, $r_{1}=1/7$, $\pi_{1}=\pi_{2}=0.1$, $\lambda=0.6$ and $\mu=0$.
@@ II.1.2 Limit of fully interdependent bees
Although all these solutions are physically valid, only two of them (one in the case $\lambda=1$) are stable. Using linear stability analysis (LSA), we determine that the stability threshold between these two solutions is determined by the highest site quality, denoted as site $k$. When $\lambda<r_{k}$, the stable solution is the absorbing state $f_{0}^{*}=1$. Thus, in the absorbing state, the system remains fully uncommitted. Conversely, when $\lambda>r_{k}$, the stable solution is the one with the best site taking the lead, with $f_{\alpha}^{*}=(1-\frac{r_{k}}{\lambda})\delta_{\alpha,k}$, and the remaining proportion of uncommitted population being $f_{0}^{*}=\frac{r_{k}}{\lambda}$. In the special case where we set $\lambda=1$, the stability threshold vanishes because $\lambda=1\geq r_{k}$, and the only stable solution remaining is $f_{0}^{*}=r_{k}$, $f_{\alpha}^{*}=(1-r_{k})\delta_{\alpha,k}$. Details on the LSA analysis are provided in Sec. A.2.
In Fig. 2B,D, we present a summary of the findings for the binary problem with $k=2$. When $\lambda\geq r_{2}$, only the fixed point where the best option is imposed remains stable, while the one where the inferior option is imposed acts as a saddle node (stable on the simplex $f_{2}=0$). As $\lambda$ decreases, these points gradually move upward (reducing either $f_{1}$ or $f_{2}$) until they merge with the point $f_{0}=1$ at $\lambda=r_{\alpha}$. In the regime where $\lambda<r_{2}$, the only remaining stable fixed point is the absorbing state $f_{0}^{*}=1$ fixed point.
@@ II.1.3 Limit of equal-quality sites
Figure 3: Color-maps illustrating the stationary values of $f_{0}$, $f_{1}$, and $f_{2}$ in the parameter space $(\pi_{1,2},\lambda)$ are presented in panels A, B, and C, respectively. The site qualities are held constant at $(q_{1},q_{2})=(7,10)$. Color gradients are used to indicate the quantitative values of $f_{\alpha}$ based on the color bars, while continuous lines represent iso-population levels in the parametric space.
@@ III Results
We will first concentrate on the specific case of a binary decision problem, a topic extensively studied in various general opinion dynamics models [11, 12, 10, 37, 38, 14, 39, 40, 41, 17]. This scenario has also been the focus of other honeybee-inspired models [22, 23, 24, 25]. The generalization to larger values of $k$ is discussed later in Section III.6.
Figure 4: A,B: Color-maps illustrating the consensus parameters $Q_{2}$ and $Q_{2}^{\prime}$, as defined in Eqs. (11) and (12) respectively with $x=2$, across the parameter space $(\pi_{1,2},\lambda$). The black lines represent the crossover points $Q_{2}=0$ or $Q_{2}^{\prime}=0$. C: Time to reach the stationary state, $t_{ss}$, obtained from simulations with a swarm size of $N=5000$ bees. The fully interdependent limit ($\pi_{\alpha}=0$ or $\lambda=1$) is not included in the maps (see discussion in Sec.III.4). The site quality parameters are held constant at $(q_{1},q_{2})=(7,10)$ across all three maps. D,E: Crossover lines $\lambda^{*}(Q_{2},Q_{2}^{\prime}=0)$ as a function of $\pi_{1,2}$. Each curve represents a different value of $q_{1}$, while $q_{2}=10$ is held constant. F: Time to reach the stationary state, $t_{ss}$, obtained from simulations plotted against the quality of the inferior option, $q_{1}$, while maintaining $q_{2}=10$, for three values of the interdependence. Parameters $\pi_{1}=\pi_{2}=0.1$ and $N=5000$ are fixed. Simulations start from uncommitted initial conditions. The markers are color-coded based on the consensus established at the stationary state. They are slightly jittered around the $q_{1}$ axis to allow for better visualization of the color code.
As detailed in the model description section, we consistently assume that site 2 holds the highest quality, denoted by $q_{2}>q_{1}$. Additionally, we assume that bees independently assess the qualities of the sites, represented by the model parameter $\mu=0$. This implies that the abandonment rates of the dances depend solely on the qualities of each site, defined as $r_{\alpha}=q_{0}/q_{\alpha}$. Previous research on this model has shown that, under these conditions, the swarm can identify the best site across a wide range of interdependence values, $\lambda$ [21, 31]. In this setup, we investigate the interplay between group communication, represented by the model parameter $\lambda$, and individual exploration success, captured by $\pi_{\alpha}$. We explore scenarios where sites may vary in their likelihood of being discovered, denoted by $\pi_{1}\neq\pi_{2}$ (asymmetric scenario), or where they share the same discovery probability, denoted by $\pi_{1}=\pi_{2}$ (symmetric scenario).
@@ III.1 Symmetric discovery scenario
Fig. 4D illustrates the consensus crossover lines for different choices of $q_{1}$, while keeping $q_{2}=10$ fixed. When we increase the quality of the bad option, consequently increasing its advertisement time, the system requires a higher degree of interdependence to favor the higher quality option. This trend would reach a turning point when $q_{1}=q_{2}$. In this scenario, the system reaches a deadlock, where $f_{1}=f_{2}$. Conversely, when $q_{1}$ is reduced, the crossover line shifts to lower values of $\lambda^{*}$, potentially leading to a situation where interdependence is not necessary to achieve consensus. For instance, the crossover line for $q_{1}=5$ lies along the x-axis, with $\lambda^{*}=0$ for all $\pi_{1,2}$. In this case, the difference in advertisement time is sufficient to maintain a significant difference between $f_{2}$ and $f_{1}$, facilitating consensus even in the absence of interdependence. Furthermore, it’s notable that the consensus crossover lines terminate abruptly at the point $(\pi_{1,2},\lambda)=(0.0,0.1)$. As we demonstrated earlier, in the fully interdependent limit ($\pi_{1,2}=0$) the system stays fully uncommitted until the threshold $\lambda=r_{2}=0.1$. Beyond this threshold, $f_{2}$ begins to take positive values while $f_{1}$ remains zero, resulting in trivial consensus.
Figure 5: Color-maps illustrating the stationary values of $f_{0}$, $f_{1}$ and $f_{2}$ are presented in panels A, B, and C, respectively, for the asymmetric discovery scenario in the parameter space $(\pi_{2}/\pi_{1},\lambda)$. The site site qualities are fixed at $(q_{1},q_{2})=(7,10)$, and the discovery probability for site 1 is set to $\pi_{1}=0.25$, while $\pi_{2}$ is changed. Color gradients indicate the quantitative values of $f_{\alpha}$ according to the color bars, while continuous lines represent iso-population levels in the parametric space.
@@
Refer to captionRefer to caption
Figure 6: A: Color-map illustrating the stationary value of the consensus parameter $Q_{2}$ (as defined from Eq. (11) with $x=2$) in the parametric space $(\pi_{2}/\pi_{1},\lambda)$. Site qualities are fixed at $(q_{1},q_{2})=(7,10)$, and the discovery probability for site 1 is set to $\pi_{1}=0.25$. The solid line indicates $Q_{2}=0$, while the dashed line indicates a simple majority, $Q_{1}=0$. B: Crossover lines $\lambda^{*}$ above which the consensus $Q_{2}$ becomes positive as the self-discovery probability $\pi_{2}$ is varied. Each curve represents a different value of site-1 quality $q_{1}$, while $q_{2}=10$ and $\pi_{1}=0.25$ are held constant. C: Crossover lines $\lambda^{*}$ above which $Q_{2}$ becomes positive as $\pi_{2}$ is varied. In this panel, qualities are held constant, $(q_{1},q_{2})=(7,10)$, and we illustrate the effect of different values of $\pi_{1}$. D: Time to reach the stationary state $t_{ss}$ in the state space ($\pi_{2}/\pi_{1},\lambda)$ obtained from simulations. Parameters $\pi_{1}=0.25$, $(q_{1},q_{2})=(7,10)$ and $N=5000$ are fixed. E: Time to reach the stationary state, $t_{ss}$, obtained from simulations plotted against the quality of the inferior option, $q_{1}$, while maintaining $q_{2}=10$, for three values of the interdependence. Parameters $(\pi_{1},\pi_{2})=(0.25,0.05)$ and $N=5000$ are fixed. Simulations start from uncommitted initial conditions. Color markers indicate the consensus value attained at the stationary state. They are slightly jittered around the $q_{1}$ axis to allow for better visualization of the color code.
@@
The trend observed is similar to that of consensus: interdependence tends to increase the decision time across all values of $\pi_{1,2}$, while the discovery probabilities speed up the process, reducing $t_{ss}$. This correlation between the established consensus value and the time to reach the stationary state suggests a speed-accuracy trade-off, commonly observed in biological systems [42] an in models of collective decision making [27, 41]. Achieving the best decision typically takes longer at higher values of $\lambda$, while processes are quicker below the consensus crossover line, where a less accurate consensus, such as a simple majority, can be reached much faster. Similarly, increasing the value of the discovery probabilities, and thus the introduction of independent information, accelerates the decision process at the cost of a lower value of $Q^{\prime}$, which can be even negative, depending on $\lambda$. Only at the limits $\lambda\rightarrow 0$ or $\lambda\rightarrow 1$ $\pi_{1,2}$ has practically no effect: in the former, there’s no discussion mediated by interdependence, resulting in a quickly established yet inaccurate stationary state, while in the latter, social feedback dominates the process, with little impact from individual exploration. It’s interesting to note the agreement between the stationary times discussed here and the relaxation times obtained from the linear stability analysis eigenvalues, as depicted in Fig. A1.